OpenStudy (anonymous):

Explain why (–4x)^0 = 1, but –4x^0 = –4

OpenStudy (irishboy123):

\[ (–4x)^0 = (–4)^0x^0 = 1 (1) = 1\] \[–4x^0 = (-4)^1 x^0 = (-4) (1) = -4\]

OpenStudy (irishboy123):

but not if x = 0!!!!